Thank you so much for your help. My problems is that Johnny wants to invest using $16000 that he has saved. The investment grew up to $25000 in 2 years.
Can you find the annual interest rate of his return by solving the following equation for him:
16000(1+x)^2 = 25000
I believe the answer is 25%, but I'm not sure how to solve equation in reverse. Can you please help me?

Answer by bacos(9)

Simplifying
16000(1 + x) * 2 = 25000
Reorder the terms for easier multiplication:
16000 * 2(1 + x) = 25000
Multiply 16000 * 2
32000(1 + x) = 25000
(1 * 32000 + x * 32000) = 25000
(32000 + 32000x) = 25000
Solving
32000 + 32000x = 25000
Solving for variable 'x'.
Move all terms containing x to the left, all other terms to the right.
Add '-32000' to each side of the equation.
32000 + -32000 + 32000x = 25000 + -32000
Combine like terms: 32000 + -32000 = 0
0 + 32000x = 25000 + -32000
32000x = 25000 + -32000
Combine like terms: 25000 + -32000 = -7000
32000x = -7000
Divide each side by '32000'.
x = -0.21875
Simplifying
x = -0.21875